OpenStudy (dhanarys):

A tree cast a shadow 10 ft long. A Boy standing next to the tree cast a shadow 2.5 ft. long. The triangle show for the tree and its shadow is similar to the triangle shadow for the boy and his shadow. If the boy is 5 ft. tall how tall is the tree?

alones (alones):

Hrmm I'm not sure but for some reason i keep on getting \(\ 20\) as my answer.

alones (alones):

Btw you have options?

OpenStudy (dhanarys):

how do you get to the answer

OpenStudy (holsteremission):

Diagram:  Because the triangles formed by the boy/his shadow and the tree/its shadow, the corresponding lengths/heights are proportional. This means \[\frac{\text{height of tree}}{\text{height of boy}}=\frac{\text{length of tree's shadow}}{\text{length of boy's shadow}}\iff\frac{x}{5}=\frac{10}{2.5}\]